need help with this problem.
x-y= 3(1)
3x-3z= 15(2)
3y+z= 6(3)

Answer by stanbon(75887)

x-y= 3
3x-3z= 15
3y+z= 6
----
Solve for "x" in 2nd equation:
x = z+5
Solve for "y" in 3rd equation:
y = (-z/3) + 2
----
Substitute into the 1st equation and solve for "z"
(z+5)-[(-z/3)+2] = 3
(4/3)z +5 - 2 = 3
(4/3)z = 0
z = 0
---
Solve for x and for "y":
x = z+5
x = 5
----
y = (-z/3)+2
y = 2
=================
Cheers,
Stan H.
============
---